OpenStudy (anonymous):

Simplify the following expression. (10x - 12y) + (49x + 67y)

whitemonsterbunny17 (whitemonsterbunny17):

Just simply add the number of x's and y's together. What do you have when you do that?

OpenStudy (anonymous):

59x + 79y -2x + 55y -2x + 79y 59x + 55y

OpenStudy (emmatassone):

\[(10x-12y)+(49x+67y)= (10x+49x)+(67y-12y)= x(10+49)+ y(67-12)= 59x+55y\]

whitemonsterbunny17 (whitemonsterbunny17):

Add the number of x's (10 + 49) to simplify the number of x's in the equation. Do the same for the number of y's (12 + 67). How many x's and y's do you have?

OpenStudy (chosenmatt):

yup

OpenStudy (anonymous):

I don't get it I try using a helpful calculator but its wrong?

whitemonsterbunny17 (whitemonsterbunny17):

Or maybe, if that doesn't make sense, I'll use colors. \(\bf\color{black}{(}\color{cyan}{10x}\color{black}{-}\color{magenta}{12y}\color{black}{)}\) + \(\bf\color{black}{(}\color{cyan}{49x}\color{black}{+}\color{magenta}{67y}\color{black}{)}\) Add the blue ones together: \(\bf\color{cyan}{10x}\color{black}{+}\color{cyan}{49x}\) to get \(\bf\color{cyan}{59x}\) Then add the y's: \(\bf\color{magenta}{12y}\color{black}{+}\color{magenta}{67y}\) to get \(\bf\color{magenta}{79y}\). So, you have \(\bf\color{cyan}{59x}\) and \(\bf\color{magenta}{79y}\) Does that make sense?

OpenStudy (anonymous):

yes thank you you get a meadle

whitemonsterbunny17 (whitemonsterbunny17):

\(\large\frak\color{blueviolet}{You're~Welcome!}\) Do you need help with anything else?

OpenStudy (anonymous):

yea with pronouns

whitemonsterbunny17 (whitemonsterbunny17):

Sure, I can help with that! Just post it in the English section and I will help. ^.^

whitemonsterbunny17 (whitemonsterbunny17):

Don't forget to close this question first, though.

OpenStudy (sleepyjess):

Bunny, it should be 67y-12y

OpenStudy (sleepyjess):

@Whitemonsterbunny17

whitemonsterbunny17 (whitemonsterbunny17):

No it shouldn't, that's not even one of the choices. ;p

OpenStudy (sleepyjess):

The answer should be 59x+55y because in the first parenthesis it has -12y

whitemonsterbunny17 (whitemonsterbunny17):

It just wanted the expression to be simplified, not the answer. :p

OpenStudy (sleepyjess):

Yes, I know but you added the y's incorrectly lol

OpenStudy (sleepyjess):

\(\bf\color{black}{(}\color{cyan}{10x}\color{red}{-}\color{magenta}{12y}\color{black}{)+}\bf\color{black}{(}\color{cyan}{49x}\color{red}{+}\color{magenta}{67y}\color{black}{)}\) One of the y values is + and one is -

whitemonsterbunny17 (whitemonsterbunny17):

Ah, I see it now. Thanks for the correction.

OpenStudy (sleepyjess):

lol we are just human XD